Someone goes skiing. He starts from rest at the top of a ski hill with an incline of 22 degrees up from the ground. If this person is going 26 m/s after 10 seconds, what is the coefficient of friction between his skis and the hill?


1
Expert's answer
2021-11-25T11:51:12-0500

"F=ma\to mg\sin22°-\mu mg\cos22°=ma\to"


"a=g\sin22°-\mu g\cos22°"


"v=at\to v=g(\sin22°-\mu \cos22°)t\to\mu=\tan22°-\frac{v}{gt\cos22°}="


"=\tan22°-\frac{26}{9.8\cdot 10\cdot\cos22°}=0.404-0.286=0.118" . Answer